# Styx: Transactional Stateful Functions on Streaming Dataflows

### v2

@@ Abstract.
Developing stateful cloud applications, such as high-throughput/ low-latency workflows and microservices with strict consistency requirements, remains arduous for programmers.
The Stateful-Functions-as-a-Service (SFaaS) paradigm aims to serve these use cases. However, existing approaches either provide serializable transactional guarantees at the level of individual functions or separate application logic from the state and use inefficient transactional protocols. These design choices increase the execution latency, limiting the usability of SFaaS systems for stateful cloud applications.
In this paper, we present Styx, a novel SFaaS runtime that executes serializable transactions across functions with exactly-once guarantees. Styx is the first streaming dataflow-based runtime for SFaaS, offering application logic and state co-location, coarse-grained state persistence, and incremental checkpointing. Styx extends a deterministic transactional protocol to support an arbitrary call graph of stateful functions. It introduces a transaction-execution acknowledgment scheme that allows tracking a transactional workflow’s SFaaS calls, guaranteeing atomicity and exactly-once processing. Experiments with the YCSB-T, TPC-C, and Deathstar benchmarks show that Styx outperforms state-of-the-art approaches by achieving at least one order of magnitude higher throughput while exhibiting near-linear scalability.
PVLDB Reference Format:
Kyriakos Psarakis, George Siachamis, George Christodoulou, Marios Fragkoulis, Asterios Katsifodimos. PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX
††This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/delftdata/styx.
@@ 1. Introduction
Despite the commercial offerings of the Functions-as-a-Service (FaaS) cloud service model, its suitability for low-latency stateful applications with strict consistency requirements, such as payment processing, reservation systems, inventory keeping, and low-latency business workflows, is quite limited. The reason behind this unsuitability is that current FaaS solutions are stateless, relying on external, fault-tolerant data stores (blob stores or databases) for state management. In addition, while multiple frameworks can perform workflow execution (e.g., AWS Step Functions (Wittig and Wittig, 2023), Azure Logic Apps), primitives for transactional execution of such applications are non-existent, contrary to the interest of users (Laigner et al., 2021; Chaoyi et al., 2023).
@@
We argue that for FaaS offerings to become mainstream, they should include state management support for stateful functions according to the Stateful-Functions-as-a-Service (SFaaS) paradigm. Overall, a suitable offering for executing workflows of stateful functions would require low-latency and high-throughput execution, end-to-end serializable transactional guarantees across multiple functions, and high-level programming models, devoid of low-level primitives for locking and transaction coordination. To the best of our knowledge, no existing approach addresses all these requirements together.
@@
The state-of-the-art transactional SFaaS with serializability guarantees Beldi (Zhang et al., 2020), Boki (Jia and Witchel, 2021a), and T-Statefun (de Heus et al., 2021a), do support transactional end-to-end workflows but suffer from high commit latency, low throughput and inadequate programming models that leak transactional primitives to developers. The main reason behind their inefficiency is the separation of state storage and function logic, as well as the use of locking and Two-Phase Commit (2PC) (GRAY, 1978) to coordinate and ensure atomicity of cross-function transactions.
This paper proposes Styx, a novel runtime for Stateful FaaS (SFaaS) that ensures exactly-once execution, low-latency, and high-throughput, while being able to execute arbitrary workflows of functions with end-to-end serializability guarantees. The guarantees offered by Styx enable the design of high-level programming models for cloud applications that obviate the need for failure management code associated with distributed systems and transaction processing inside application logic. In Styx, transactions are handled with a deterministic transaction protocol, avoiding costly 2PCs. As shown in Figure 1, Styx achieves at least one order of magnitude lower median latency, two orders of magnitude lower 99p latency at 2000 transactions per second, and at least one order of magnitude higher throughput compared to the state of the art SFaaS systems (Jia and Witchel, 2021a; Zhang et al., 2020; de Heus et al., 2021a) with end-to-end serializable guarantees.
In short, this paper makes the following contributions.
Styx is a scalable, transactional, and fault-tolerant SFaaS runtime that leverages the streaming dataflow execution model and co-locates execution with state to offer exactly-once processing guarantees (Section 4).
@@
Styx provides multi-partition transactional serializability across arbitrary function calls (Section 5 and Section 6), extending the concept of deterministic databases for stateful functions and contributing a novel acknowledgment-sharing scheme (Section 5.3) to track function execution efficiently.
Styx contributes a high-level programming model for transactional stateful functions, abstracting away transaction and failure management code (Section 3).
@@
Styx’s asynchronous incremental snapshots achieve fault-tolerance with minimal overhead to the system’s operation (Section 7), while Styx’s determinism enables early commit replies: the means to enable transactions to be reported as committed before a snapshot is committed to durable storage (Section 6.4).
Outline. The rest of the paper is organized as follows. Section 2 provides the background and highlights the motivation behind our work. Then, in Section 3, we present Styx’s programming model. In Section 4, we analyze the architecture of Styx. We delve into the sequencing & function execution workflow in Section 5, followed by Section 6, which discusses how we handle the deterministic commit of transactions. We end the technical description of Styx in Section 7 with the description of the fault tolerance mechanisms based on incremental snapshots. We devote Section 8 to our experimental methodology, results, and observations. Finally, we discuss the related work in Section 9 and conclude our paper in Section 10.
We make Styx publicly available: https://github.com/delftdata/styx
2. Background and Motivation
Although relational database systems provide ACID transaction guarantees, many distributed real-world applications suffer from serious consistency issues when the responsibility of transaction execution is not ensured by a database system and is left to developers (Chaoyi et al., 2023; Laigner et al., 2021).
In this section, we delve into the specifics of streaming dataflow systems design and argue that they can be extended to encapsulate the primitives required for executing stateful functions consistently and efficiently. We then show how the concept of deterministic databases can be extended for SFaaS, where transaction boundaries are unknown, unlike online transaction processing (OLTP). Finally, we argue that combining deterministic databases and dataflow systems can create a runtime that ensures atomicity, consistency, and scalability.
@@ 2.1. Dataflows for Stateful Functions
Stateful dataflows is the underlying execution model implemented by virtually all modern stream processors (Carbone et al., 2015; Murray et al., 2013; Noghabi et al., 2017). Besides being a great fit for parallel, data-intensive computations, stateful dataflows are the primary abstraction supporting workflow managers such as Apache Airflow (Harenslak and de Ruiter, 2021), AWS Step Functions (Wittig and Wittig, 2023), and Azure’s Durable Functions (Burckhardt et al., 2021a). We argue that stateful dataflows can make a suitable runtime for orchestrating general-purpose cloud applications for the following reasons.
Exactly-once Processing. Message-delivery guarantees are fundamentally hard to deal with in the general case, with the root of the problem being the well-known Byzantine Generals problem (Lamport et al., 1982). However, in the closed world of dataflow systems, exactly-once processing is possible (Carbone et al., 2015, 2017; Silvestre et al., 2021). As a matter of fact, the APIs of popular streaming dataflow systems, such as Apache Flink, require no error management code whatsoever (e.g., message retries or duplicate elimination with idempotency IDs).
Co-location of State and Function. The primary reason dataflow systems can sustain millions of events per second (Carbone et al., 2015; Gencer et al., 2021) is that their state is partitioned across operators that operate on local state. While this co-location goes against the current economics of Cloud offerings that want state and computation to be decoupled, we believe that it is the primary vehicle to high-performance, and should be adopted by modern SFaaS runtimes.
@@ 2.2. Determinism & Transactions
Deterministic databases advance the level of transactional guarantees provided by traditional database systems. Given a set of database partitions and a set of transactions, a deterministic database (Abadi and Faleiro, 2018; Thomson and Abadi, 2010) will end up in the same final state despite node failures and possible concurrency issues.
@@
Traditional database systems offer serializable guarantees, which allow multiple transactions to execute concurrently, ensuring that the database state will be equivalent to a serial execution of the transactions. Thus, deterministic databases guarantee not only serializability but also that a given set of transactions will have exactly the same effect on the database state despite transaction re-execution. This guarantee has multiple important implications (Abadi and Faleiro, 2018) that, to the best of our knowledge, have not been leveraged by SFaaS systems thus far.
@@
Deterministic Databases on Streaming Dataflows. Unlike 2PC, which requires rollbacks in case of failures, deterministic database protocols (Lu et al., 2020; Thomson et al., 2012) are ”forward-only”: once the locking order (Thomson et al., 2012) or read/write set (Lu et al., 2020) of a batch of transactions has been determined, the transactions are going to be executed and reflected on the database state, without the need to rollback changes. This is very much in line with how dataflow systems operate: events flow through the dataflow graph, from sources to sinks, without stalls for coordination. This match between deterministic databases and the dataflow execution model is the primary motivation behind Styx’s design choice to implement a deterministic transaction protocol on top of a dataflow system.
Non-Deterministic Functions on Streaming Dataflows. In principle, functions may encapsulate logic that makes the outcome of their execution non-deterministic. Examples of non-deterministic operations are calls to external systems and the use of random number generators or time-related utilities. Styx, like state-of-the-art SFaaS systems (Zhang et al., 2020; Jia and Witchel, 2021a), currently supports functions with deterministic logic that always leads to the same output given the same input. That said, it is possible to embrace non-deterministic functions in Styx by tracking, recording, and persisting non-deterministic logic contained in them following the approach of Clonos (Silvestre et al., 2021).
2.3. Challenges & Open Problems
@@
Despite their success and widespread applicability, dataflow systems need to undergo multiple changes before they can be used for transactional stateful functions. In the following, we list the challenges and open problems we tackle in this paper.
@@
Programming Models. Dataflow systems at the moment are only programmable through functional-programming style dataflow APIs: a given cloud application has to be rewritten by programmers to match the event-driven dataflow paradigm. Although it is possible to rewrite many applications in this paradigm, it takes a considerable amount of programmer training and effort to do so. We argue that dataflow systems would benefit from object-oriented programming abstractions in order to be adopted by programmers for general cloud applications, such as microservices.
Support for Transactions. Although it has been recently shown that it is possible to introduce transaction coordination on top of a dataflow system (de Heus et al., 2021a), we argue that a more efficient implementation of transactions would require the dataflow system to be aware of transaction boundaries and to incorporate transaction processing into its state management and fault-tolerance protocols.
Deterministic OLTP and SFaaS. OLTP databases that use deterministic protocols like Calvin (Thomson et al., 2012; Zhou et al., 2022) require each transaction’s read/write set a priori. Aria (Lu et al., 2020) extended deterministic transaction protocols by discovering the read-write sets of a transaction by first executing it. Aria works under the assumption that a transaction is encapsulated in a single-threaded function that can execute remote reads and writes from multiple partitions.
@@
Arbitrary function calls enable programmers to take advantage of both the separation of concerns principle that is widely applied in microservice architectures (Laigner et al., 2021), as well as code modularity. Although deterministic database systems have been proven to perform exceptionally well (Abadi and Faleiro, 2018), the design and implementation of a deterministic transactional protocol for a SFaaS runtime with arbitrary function call graphs is non-trivial. More specifically, arbitrary function calls create complex call graphs that need to be tracked in order to establish a transaction’s boundaries before committing. At the same time, current deterministic transaction approaches employ a single thread that performs remote reads and writes instead of making arbitrary calls in parallel to functions that can mutate state.
@@
Dataflows for Arbitrary-Workflow Execution. The prime use case for dataflow systems nowadays is streaming analytics. However, general cloud applications have different workload characteristics. Functions calling other functions and receiving return values introduce cycles in the dataflow graph. Such cycles can cause deadlocks and other flow-control issues that need to be dealt with (Lattuada et al., 2016).
@@
In this work, we tackle these challenges and propose Styx: a novel dataflow system tailored to the needs of stateful functions, with built-in support for deterministic transactions and a high-level programming model.
@@ 3. Programming Model
The programming model of Styx is based on Python and comprises operators that encapsulate partitioned mutable state and functions that operate on that. An example showcasing the programming model of Styx can be found in Figure 2.
@@ 3.1. Programming Model Notions
Stateful Entities. Similar to objects in object-oriented programming, entities in Styx are responsible for mutating their own state. Moreover, when a given entity needs to update the state of another entity, it can do so via a function call. Each entity bears a unique and immutable key, similar to Actor references in Akka (Akk, [n.d.]), with the difference that entity keys are application-dependent and contain no information related to their physical location. The dataflow runtime engine (Section 4) uses that key to route function calls to the right operator that accommodates that specific entity.
@@
Functions. functions can mutate the state of an entity. By convention, the context is the first parameter of each function call. Functions are allowed to call other functions directly, and Styx supports both synchronous and asynchronous function calls. For instance, in lines 9-11 of Figure 2, the instantiated reservation entity will call asynchronously the function ’reserve_hotel’ of an entity with key ’hotel_id’ attached to the Hotel operator. Similarly, one can make a synchronous call that blocks to wait for results. In this case, Styx will block execution until the call returns. Depending on the use case, a mix of synchronous and asynchronous calls can be used. Asynchronous function calls, however, allow for further optimizations that Styx applies whenever possible, as we describe in Section 5 and Section 6.
@@
Operators. Each entity directly maps to a dataflow operator (also called a vertex) in the dataflow graph. When an event (i.e., function invocation) enters the dataflow graph, it reaches the operator holding the function code of the given entity as well as the state of that entity. In short, a dataflow operator can execute all functions of a given entity and store the state of that entity. Since operators can be partitioned across multiple cluster nodes, each partition stores a set of stateful entities indexed by their unique key. When a function of an entity is invoked (via an incoming event), the entity’s state is retrieved from the local operator state. Then, the function is executed using the arguments found in the incoming event that triggered the call.
@@
State & Namespacing. As mentioned before, each entity has access only to its own state. In Styx, the state is namespaced with respect to the entity it belongs to. For instance, a given key ”hotel53” within the operator Hotel is represented as: entities://Hotel/hotel53. This way, a reference to a given key of a state object is unique and can be determined at runtime when operators are partitioned across workers. Programmers can store or retrieve state through the context object by invoking context.put() or get() (e.g., in 17 of Figure 2). Styx’s context is similar to the context object used in other systems such as Flink Statefun, AWS Lambda, and Azure Durable Functions.
@@
Transactions. A transaction in Styx begins with the invocation of an ”entry” function. The subsequent functions that are called from the entry function (call graph) form a transaction that executes with serializable guarantees. Styx’s programming model allows transaction aborts by raising an uncaught exception. In the example of Figure 2, if a hotel entity does not have enough availability when calling the ’reserve_hotel’ function, the ’make_reservation’ transaction should be aborted, alongside potential state mutations that the ’reserve_flight’ has made to a flight entity. In that case, the programmer has to raise an exception as follows:
@@
Exactly-once Function Calling. A significant contribution of Styx is that it offers exactly-once processing guarantees: Styx reflects the state changes of a function call execution exactly-once. This way, programmers do not need to “pollute” their business logic with consistency checks, state rollbacks, timeouts, retries, and idempotency (Laigner et al., 2021; Killalea, 2016). We detail this capability in Section 7.

### v4

@@ Abstract.
Developing stateful cloud applications, such as low-latency workflows and microservices with strict consistency requirements, remains arduous for programmers. The Stateful Functions-as-a-Service (SFaaS) paradigm aims to serve these use cases. However, existing approaches provide weak transactional guarantees or perform expensive external state accesses requiring inefficient transactional protocols that increase execution latency.
In this paper, we present Styx, a novel dataflow-based SFaaS runtime that executes serializable transactions consisting of stateful functions that form arbitrary call-graphs with exactly-once guarantees. Styx extends a deterministic transactional protocol by contributing: i) a function acknowledgment scheme to determine transaction boundaries required in SFaaS workloads, ii) a function-execution caching mechanism, and iii) an early-commit reply mechanism that substantially reduces transaction execution latency. Experiments with the YCSB, TPC-C, and Deathstar benchmarks show that Styx outperforms state-of-the-art approaches by achieving at least one order of magnitude higher throughput while exhibiting near-linear scalability and low latency.
@@ 1. Introduction
Despite the commercial offerings of the Functions-as-a-Service (FaaS) cloud service model, its suitability for low-latency stateful applications with strict consistency requirements, such as payment processing, reservation systems, inventory keeping, and low-latency business workflows, is quite limited. The reason behind this unsuitability is that current FaaS solutions are stateless, relying on external, fault-tolerant data stores (blob stores or databases) for state management. In addition, while multiple frameworks can perform workflow execution (e.g., AWS Step Functions (Wittig and Wittig, 2023), Azure Logic Apps (Corporation, 2025)), they do not provide primitives for transactional execution of such applications. As a result, distributed applications (e.g., microservice architectures) suffer from serious consistency issues when the responsibility of transaction execution is left to developers (Chaoyi et al., 2023; Laigner et al., 2021).
@@
In line with recent research (Sreekanti et al., 2020; Jia and Witchel, 2021a; Zhang et al., 2020; de Heus et al., 2021; Jia and Witchel, 2021b; Spenger et al., 2022), we agree that for FaaS offerings to become mainstream, they should include state management support for stateful functions according to the Stateful Functions-as-a-Service (SFaaS) paradigm. In addition, we argue that a suitable runtime for executing workflows of stateful functions should also provide end-to-end serializable transactional guarantees across multiple functions, low-latency and high-throughput execution, and a high-level programming model, devoid of low-level primitives for locking and transaction coordination. To the best of our knowledge, no existing approach addresses all these requirements together.
@@
The state-of-the-art transactional SFaaS with serializable guarantees, Boki (Jia and Witchel, 2021a), Beldi (Zhang et al., 2020), and T-Statefun (de Heus et al., 2021) do support transactional end-to-end workflows but induce high commit latency and low throughput. The main reason behind their inefficiency is the separation of state storage and function logic, as well as the use of locking and Two-Phase Commit (2PC) (GRAY, 1978) to coordinate and ensure the atomicity of cross-function transactions.
This paper proposes Styx, a novel dataflow-based runtime for SFaaS. Styx ensures that each transaction’s state mutations will be reflected once in the system’s state, even under failures, retries, or other potential disruptions (known as exactly-once processing). Additionally, Styx can execute arbitrary function orchestrations with end-to-end serializability guarantees, leveraging concepts from deterministic databases to avoid costly 2PCs.
Our work stems from two important observations. First, modern streaming dataflow systems such as Apache Flink (Carbone et al., 2015) guarantee exactly-once processing (Carbone et al., 2015, 2017; Silvestre et al., 2021) by transparently handling failures. A limitation of those streaming systems is that they cannot execute general cloud applications such as microservices or guarantee transactional SFaaS orchestrations. Second, deterministic database protocols (Thomson et al., 2012; Lu et al., 2020) that can avoid expensive 2PC invocations have not been designed for complex function orchestrations and arbitrary call-graphs. For the needs of transactional SFaaS, Styx leverages a deterministic transactional protocol, enabling early commit replies to clients (i.e., before a snapshot is committed to persistent storage).
Our work is in line with recent proposals in the area, such as DBOS (Skiadopoulos et al., 2021), Hydro (Cheung et al., 2021), and SSMSs (Li et al., 2024). Contrary to these systems, our work adopts the streaming dataflow execution model and guarantees serializability across functions. As shown in Figure 1, Styx achieves one order of magnitude lower median latency, two orders of magnitude lower 99p latency at 2000 transactions/sec, and one order of magnitude higher throughput compared to state of the art (SotA) serializable SFaaS systems (Jia and Witchel, 2021a; Zhang et al., 2020; de Heus et al., 2021).
In short, this paper makes the following contributions:
– Styx combines deterministic transactions with dataflows and overcomes the challenges that arise from this design choice (Section 2).
– Styx enables high-level SFaaS programming models that abstract away transaction and failure management code (Section 3). Styx does so, by guaranteeing exactly-once processing (Section 4) and transactional serializability across arbitrary function calls (Section 5 and Section 6).
@@
– Styx extends the concept of deterministic databases to support arbitrary workflows of stateful functions, contributing a novel acknowledgment scheme (Section 5.3) to track function completion efficiently, as well as a function-execution caching mechanism (Section 6.3) to speed up function re-executions.
@@
– Styx’s deterministic execution enables early commit replies: transactions can be reported as committed, even before a snapshot of executed transactions is committed to durable storage (Section 6.4).
– Styx outperforms the state-of-the-art (Zhang et al., 2020; Jia and Witchel, 2021a; de Heus et al., 2021) by at least one order of magnitude higher throughput in all tested workloads while achieving lower latency and near-linear scalability (Section 8).
Styx is available at: https://github.com/delftdata/styx
2. Motivation
In this section, we analyze the specifics of streaming dataflow systems design and argue that they can be extended to encapsulate the primitives required for consistently and efficiently executing workflows of stateful functions. Our work is based on a key observation: the architecture of high-performance cloud services closely resembles a parallel dataflow graph, where the state is partitioned and co-located with the application logic (Psarakis et al., 2025). Additionally, as we detail in Section 2.2, there is a synergy between deterministic transactions and dataflow systems. Such a combination can offer state consistency and ease of programming as monolithic solutions did in the past, while improving scalability and eliminating developer involvement. Finally, we show how deterministic transactions can be extended for SFaaS, where transaction boundaries are unknown, unlike online transaction processing (OLTP).
@@ 2.1. Dataflows for Stateful Functions
Stateful dataflows is the execution model implemented by virtually all modern stream processors (Carbone et al., 2015; Murray et al., 2013; Noghabi et al., 2017). Besides being a great fit for parallel, data-intensive computations, stateful dataflows are the primary abstraction supporting workflow managers such as Apache Airflow (Harenslak and de Ruiter, 2021), AWS Step Functions (Wittig and Wittig, 2023), and Azure’s Durable Functions (Burckhardt et al., 2021a). In the following, we present the primary motivation behind using stateful dataflows to build a suitable runtime for orchestrating general-purpose cloud applications.
Exactly-once Processing. Message-delivery guarantees are fundamentally hard to deal with in the general case, with the root of the problem being the well-known Byzantine Generals problem (Lamport et al., 1982). However, in the closed world of dataflow systems, exactly-once processing is possible (Carbone et al., 2015, 2017; Silvestre et al., 2021). As a matter of fact, the APIs of popular streaming dataflow systems, such as Apache Flink, require no error management code (e.g., message retries or duplicate elimination with idempotency IDs).
Co-Location of State and Function. The primary reason streaming dataflow systems can sustain millions of events per second (Carbone et al., 2015; Gencer et al., 2021) is that their state is partitioned across operators that operate on local state. While the structure of current Cloud offerings favors the disaggregation of storage and computation, we argue that co-locating state and computation is the primary vehicle for high performance and can also be adopted by modern SFaaS runtimes, as opposed to using external databases for state storage.
Coarse-Grained Fault Tolerance. To ensure atomicity at the level of workflow execution, existing SFaaS systems perform fine-grained fault tolerance (Zhang et al., 2020; Jia and Witchel, 2021a); each function execution is logged and persisted in a shared log before the next function is called. This requires a round-trip to the logging mechanism for each function call, which adds significant latency to function execution. Instead of logging each function execution, streaming dataflow systems (Siachamis et al., 2024; Carbone et al., 2017; Chandy and Lamport, 1985) opt for a coarse-grained fault tolerance mechanism based on asynchronous snapshots, reducing this overhead.
@@ 2.2. Determinism & Transactions
Given a set of database partitions and a set of transactions, a deterministic database (Abadi and Faleiro, 2018; Thomson and Abadi, 2010) will end up in the same final state despite node failures and possible concurrency issues. Traditional database systems offer serializable guarantees, allowing multiple transactions to execute concurrently, ensuring that the database state will be equivalent to the state of one serial transaction execution. Deterministic databases guarantee not only serializability but also that a given set of transactions will have exactly the same effect on the database state despite transaction re-execution. This guarantee has important implications (Abadi and Faleiro, 2018) that have not been leveraged by SFaaS systems thus far.
@@
Deterministic Transactions on Streaming Dataflows. Unlike 2PC, which requires rollbacks in case of failures, deterministic database protocols (Lu et al., 2020; Thomson et al., 2012) are ”forward-only”: once the locking order (Thomson et al., 2012) or read/write set (Lu et al., 2020) of a batch of transactions has been determined, the transactions are going to be executed and reflected on the database state, without the need to rollback changes. This notion is in line with how dataflow systems operate: events flow through the dataflow graph, from sources to sinks, without stalls for coordination. This match between deterministic databases and the dataflow execution model is the primary motivation behind Styx’s design choice to implement a deterministic transaction protocol on top of a dataflow system.
2.3. Challenges
@@
Despite their success and widespread applicability, dataflow systems need to undergo multiple changes before they can be used for transactional stateful functions. In the following, we list challenges and open problems tackled in this work.
@@
Programming Models. Dataflow systems at the moment are only programmable through functional programming-style dataflow APIs: a given cloud application has to be rewritten by programmers to match the event-driven dataflow paradigm. Although it is possible to rewrite many applications in this paradigm, it takes a considerable amount of programmer training and effort. We argue that dataflow systems would benefit from object-oriented or actor-like programming abstractions in order to be adopted for general cloud applications, such as microservices.
Support for Transactions. Transactions in the context of streaming dataflow systems typically refer to processing a set of input elements and their state updates with ACID guarantees (Zhang et al., 2024). Despite progress, critical challenges remain open, such as the performance overhead incurred by multi-partition transactions, as well as the need to block flows of data for locking and message re-ordering. In this work, we argue that in order to implement transactions in a streaming dataflow system, we need to ”keep the data moving” (Stonebraker et al., 2005) by avoiding disruptions in the natural flow of data while tightly integrating transaction processing into the system’s state management and fault tolerance protocols.
@@
Deterministic OLTP and SFaaS. OLTP databases that use deterministic protocols like Calvin (Thomson et al., 2012; Zhou et al., 2022; Lu et al., 2020) either require each transaction’s read/write set a priori or are extended to discover the read-write sets of a transaction by first executing it. Additionally, in both scenarios, deterministic protocols assume that a transaction is executed as a single-threaded function that can perform remote reads and writes from other partitions. In the case of SFaaS, arbitrary function calls enable programmers to take advantage of both the separation of concerns principle, which is widely applied in microservice architectures (Laigner et al., 2021), as well as code modularity. Although deterministic database systems have been proven to perform exceptionally well (Abadi and Faleiro, 2018), designing and implementing a deterministic transactional protocol for arbitrary workflows of stateful functions is non-trivial. Specifically, arbitrary function calls create complex call-graphs that need to be tracked in order to establish a transaction’s boundaries before committing.
@@
Dataflows for Arbitrary-Workflow Execution. The prime use case for dataflow systems nowadays is streaming analytics. However, general-purpose cloud applications have different workload requirements. Functions calling other functions and receiving responses introduce cycles in the dataflow graph. Such cycles can cause deadlocks and need to be dealt with (Lattuada et al., 2016).
@@
In this work, we tackle these challenges and propose a dataflow system tailored to the needs of stateful functions with built-in support for deterministic transactions and a high-level programming model.
@@ 3. Programming Model
The programming model of Styx is based on Python and comprises operators that encapsulate partitioned mutable state and functions that operate on that. An example of the programming model of Styx is depicted in Figure 2.
@@ 3.1. Programming Model Notions
Stateful Entities. Similar to objects in object-oriented programming, entities in Styx are responsible for maintaining and mutating their own state. Moreover, when a given entity needs to update the state of another entity, it can do so via a function call. Each entity bears a unique and immutable key, similar to Actor references in Akka (Akk, [n. d.]), with the difference that entity keys are application-dependent and contain no information related to their physical location. The dataflow runtime engine (Section 4) uses that key to route function calls to the right operator that accommodates that specific entity.
@@
Functions. functions can mutate the state of an entity. By convention, the context is the first parameter of each function call. Functions are allowed to call other functions directly, and Styx supports both synchronous and asynchronous function calls. For instance, in lines 9-11 of Figure 2, the instantiated reservation entity will call asynchronously the function ’reserve_hotel’ of an entity with key ’hotel_id’ attached to the Hotel operator. Similarly, one can make a synchronous call that blocks waiting for results. In this case, Styx will block execution until the call returns. Depending on the use case, a mix of synchronous and asynchronous calls can be used. Asynchronous function calls, however, allow for further optimizations that Styx applies whenever possible, as we describe in Section 5 and Section 6.
@@
Operators. Each entity directly maps to a dataflow operator (also called a vertex) in the dataflow graph. When an event enters the dataflow graph, it reaches the operator holding the function code of the given entity as well as the state of that entity. In short, a dataflow operator can execute all functions of a given entity and store the state of that entity. Since operators can be partitioned across multiple cluster nodes, each partition stores a set of stateful entities indexed by their unique key. When an entity’s function is invoked (via an incoming event), the entity’s state is retrieved from the local operator state. Then, the function is executed using the arguments found in the incoming event that triggered the call.
@@
State & Namespacing. As mentioned before, each entity has access only to its own state. In Styx, the state is namespaced with respect to the entity it belongs to. For instance, a given key ”hotel53” within the operator Hotel is represented as: entities://Hotel/hotel53. This way, a reference to a given key of a state object is unique and can be determined at runtime when operators are partitioned across workers. Programmers can store or retrieve state through the context object by invoking context.put() or get() (e.g., in Line 17 of Figure 2). Styx’s context is similar to the context object used in other systems such as Flink Statefun, AWS Lambda, and Azure Durable Functions.
@@
Transactions. A transaction in Styx begins with a client request. The functions that are part of the transaction form a workflow that executes with serializable guarantees. Styx’s programming model allows transaction aborts by raising an uncaught exception. In the example of Figure 2, if a hotel entity does not have enough availability when calling the ’reserve_hotel’ function, the ’make_reservation’ transaction should be aborted, alongside potential state mutations that the ’reserve_flight’ has made to a flight entity. In that case, the programmer has to raise an exception as follows:
@@
Exactly-once Function Calling. Styx offers exactly-once processing guarantees: it reflects the state changes of a function call execution exactly-once. Thus, programmers do not need to “pollute” their application logic with consistency checks, state rollbacks, timeouts, retries, and idempotency (Laigner et al., 2021; Killalea, 2016). We detail this capability in Section 7.